Olive wants to invest $15,000 with the goal of earning 10% interest. She wants to put some of the money in an investment that has been earning 7% a year and the rest of the money in an investment that typically earns 11% a year. How much money should be invested at 11% interest for Olive to obtain her goal of an overall 10% interest?
Answer by richwmiller(17219)

x+y=15000,
0.07*x+0.11*y=1500
x=15000-y
0.07*(15000-y)+0.11*y=1500
1050.0-0.07y+0.11*y=1500
0.04*y=450.0
y=11250.0
x=15000-y
x=3750.0 y=11250.0
check
3750.0 +11250.0=15000
0.07*x+0.11*y=1500
0.07*3750.0+0.11*11250.0=1500
262.5+1237.5=1500
1500.0=1500
ok